OpenStudy (anonymous):

What is the slope of the line passing through the points (2, 7) and (–1, 3)?

OpenStudy (hoblos):

slope = (3-7)/(-1-2) = -4/-3 = 4/3

OpenStudy (anonymous):

Thank you!

Directrix (directrix):

Slope of the line is change in y over change in x. Change in y = 7-4 = 4 Change in x = 2 - (-1) = 3 Slope: 4/3